50.0 grams of C_16H_34 are burned at 1 bar pressur Both a and b
Answer #1

a) balanced equation :

2C16H34 + 47O2 ---> 28CO2 + 4CO + 34H2O

dHrxn = -1945.2 kJ

moles of C16H34 = 50 g/226.44 g/mol = 0.22 mols

2 mole of C16H34 reacts to form 4 moles of CO

So fraction of hexadecane forming CO = 0.22 x 4/2 = 0.44 mols

b) Change in internal energy = -1945.2/0.44 = 4420.91 kJ/mol